OpenStudy (anonymous):

prove that if \( 0 < a < b \) then \[a < \sqrt{ab} < \frac{a+b}{2} < b\]

OpenStudy (anonymous):

\[(a+b)/2\] is average value of a and b so obviously it will be between a and b \[\sqrt{ab}\] since a is smaller than b \[\sqrt{ab}<\sqrt{b*b}=b\] \[\sqrt{ab}>\sqrt{a*a}=a\]

OpenStudy (anonymous):

it remains to show that \( \sqrt{ab} \leq \frac{a+b}{2},\ \ \ a,b \geq 0\)

OpenStudy (anonymous):

thinking.....

OpenStudy (anonymous):

square them both \[\frac{a^2}{4}+\frac{a b}{2}+\frac{b^2}{4}>\text{ab}\]

OpenStudy (anonymous):

equivalent to \[a^2+b^2>2ab\]

OpenStudy (anonymous):

\[a^2+b^2-2ab>0\] \[(a-b)^2>0\] which is true.

OpenStudy (anonymous):

If \(a < b \) then \[a = \frac{a+a}{2} < \frac{a+b}{2} < \frac{b+b}{2} < b\] now if \( 0 < a < b\) then \( a^2 < ab\) so \( a < \sqrt{ab}\)\[(a-b)^2 > 0 \implies \]\[a^2 + b^2 >2ab\]\[a^2 + 2ab + b^2 > 4ab\]\[(a+b)^2 > 4ab\]so\( a+b > 2 \sqrt{ab} \). Moreover, for all a,b, we have \( (a-b)^2 \geq 0 \), and thus \( (a+b)^2 \geq 4ab\) which implies that \( a+b \geq 2\sqrt{ab} \) for all \(a,b \geq 0 \)

OpenStudy (anonymous):

looks good